# Graph Neural Network-Based Track Finding in the LHCb Vertex Detector

### v2

@@ 2 The LHCb detector and computing framework
The LHCb data acquisition can sustain a peak data rate of 5 TB/s which must be reduced to 10 GB/s before being written to long-term storage. This reduction is performed by a two-stage real-time processing system. The first stage is implemented on GPUs using the Allen [^2] https://gitlab.cern.ch/lhcb/Allen [2] framework. Allen performs a partial event reconstruction, including tracking, and reduces the rate to between 70 and 200 GB/s using around 50 physics selections. The second stage is implemented on CPUs using the Moore [^3] https://gitlab.cern.ch/lhcb/Moore application based on the Gaudi [^4] https://gitlab.cern.ch/gaudi framework. Moore performs a full offline-quality reconstruction of the LHCb detector and reduces the rate to 10 GB/s using over two thousand physics selections.
Given this context, Allen and LHCb offer an interesting use-case for developing Neural Network-based algorithms on GPUs, with a specific emphasis on high-throughput performance. In particular LHCb processes the full LHC collision rate using only around 320 GPU cards, leading to a minimum required throughput of around 80 kHz for the full processing pipeline on a single GPU. This in turn makes it essential to minimize overheads caused by data transfers between the host server and the GPU, which Allen achieves by processing batches of O(1000) events at a time. To the best of our knowledge this is the first time that a GNN tracking pipeline has been implemented using batched processing of this kind, or targeting these kinds of throughputs per GPU.
3 Datasets
The results presented in this paper have been obtained using events produced with the full LHCb detector simulation in which $p$-$p$ collisions are generated using PYTHIA [22] and decayed using EvtGen [23]. The interaction of the generated particles with the detector, and its response, are implemented using the Geant4 toolkit [24] as described in [25]. We use simulated minimum bias samples corresponding to typical LHCb datataking conditions between 2022 and 2025, with an average of 5.3 inelastic $p$-$p$ collisions which produce at least one particle with momentum above 2 GeV in the LHCb detector acceptance per event. In our sample, there are, on average, 150 particles in the VELO acceptance, and 2,200 hits, in each event. In the simulation used for this paper around 15% of the hits are spilled over from prior events and therefore noise. The embedding network and GNN are trained on 700,000 events, adhering to these selection criteria:
  • The hits of the particle tracks that are not sufficiently linear due to multiple scattering (largely soft electrons) are removed. This is assessed by fitting a line to the particle hits and applying an upper limit to the average squared distance between the hits and the line. This criterion improves the physics performance during the training but also excludes 2.5% of the tracks reconstructible in the VELO.
  • A minimum of 500 genuine VELO hits is required.
  • Tracks with fewer than 3 hits are excluded.
These criteria are not imposed on the test samples. Allen’s existing tracking algorithms, which are used by LHCb in 2024 datataking, are used as a reference for both the physics and computational performance of ETX4VELO.
@@ 4 Track Topologies in the VELO
A VELO plane consists of four overlapping sensor layers, displaced in $z$, collaboratively covering the desired acceptance in the $\left(x,y\right)$. The effective hit efficiency of the VELO is around 99% in simulation, so tracks must be allowed to skip a VELO plane during the reconstruction. A track can also leave more than one hit per plane, in most cases when it traverses overlapping sensor layers within a given plane. Material interactions frequently produce positron-electron pairs, resulting in two tracks that initially share hits before diverging. In fact 55% of electrons share hits with another particle, which motivates the need to take particular care when reconstructing them. Two tracks may also accidentally intersect, leading to a shared hit. The tracks that an effective VELO tracking algorithm should be capable of reconstructing are illustrated in Figure 2.
@@ 5 The ETX4VELO Pipeline
The hits left by charged particles in the VELO detector can be thought of as a graph, in which hits originating from the same particle are connected to each other. The basic idea of a GNN pipeline is to build the graph of possible connections between hits in the detector, accurately classify hits as connected or unconnected, and then transform these connections into a track object which can be understood and used by the rest of LHCb’s real-time pipeline. Since a graph consisting of all possible hit connections would be prohibitively large, a key challenge is to construct the initial graph in such a way that nearly all connected hits are part of the graph, while as many unconnected hits as possible are not. Our pipeline consists of 7 steps, as illustrated in Figure 3. We summarize the steps here, and then give further details on each step in dedicated subsections.
Starting with the hits alone, the first two steps of the ETX4VELO pipeline construct a rough graph, $G^{\text{hit}}_{\text{rough}}$. In what follows we refer to an “edge” as a connection between two hits in the detector. Hits are first embedded into an Euclidean space using a Multi-Layer Perceptron (MLP). The MLP is trained to position hits that are likely to be connected by an edge close to each other in the embedding space. Subsequently, $k$-Nearest Neighbours ($k$-NN) algorithms are applied in the embedding space to collect edges that are most likely to be genuine. Nodes are considered at most two planes apart when identifying edge candidates.
@@
First, a GNN encodes each edge in $G^{\text{hit}}_{\text{rough}}$ ($i\to j$) into a high-dimensional vector $e_{i\to j}$, with dual training targets: edge classification and triplet classification. Second, an MLP (edge classifier) transforms these encodings into edge scores ranging from 0 (fake) to 1 (genuine). Edges with scores below $s_{\text{edge, min}}$ are discarded, resulting in the purified hit graph $G^{\text{hit}}_{\text{purified}}$. Third, the graph of edges $G^{\text{edge}}$ is built from the purified hit graph $G^{\text{hit}}_{\text{purified}}$. Finally, a second MLP (triplet classifier) transforms the pairs of edge encodings (triplets) into triplet scores. Triplets scoring below $s_{\text{edge, min}}$ are discarded, resulting in the purified edge graph $G^{\text{edge}}_{\text{purified}}$.
@@ 5.1 Hit Embedding and Rough Graph Construction
To construct the graph $G^{\text{hit}}_{\text{rough}}$, one could connect each hit to all hits on the next two planes, allowing for the possibility of a missing plane due to pixel inefficiencies. However, this approach results in an excessive number of edges, thereby increasing the GNN’s inference time and memory usage. To enhance throughput, it is essential to minimise the graph size at this stage.
@@
Using this trained embedding MLP, a hit on plane $p$ is connected to hits on the next two planes, $p+1$ and $p+2$, if they are within a squared distance of $d^{2}_{\text{max}}$. To avoid an excessive number of edges, a maximum of $k_{\text{max}}$ edges per node is imposed. Consequently, the rough graph $G^{\text{hit}}_{\text{rough}}$ is constructed by applying a $k_{\text{max}}$-NN algorithm on plane $p\in\left\llbracket 0,n_{\text{planes}}-1\right\rrbracket$ to the next two planes, $p+1$ and $p+2$, under a maximum squared distance of $d^{2}_{\text{max}}$. The k-NN implementation from faiss [28] is used for this purpose. The values $k_{\text{max}}=50$ and $d^{2}_{\text{max}}=0.9$ are determined post-training.
@@
The margin $m$, representing a squared distance threshold, is fixed at $m=1$. The parameter $w_{\text{genuine}}>1$ reduces the likelihood of excluding true edges within this margin, thus favouring the inclusion of genuine edges over the exclusion of false ones.
@@ 5.2 Graph Neural Network and Classifiers
The Graph Neural Network (GNN) is employed to derive edge encodings used for both edge and triplet classification. The architecture of the GNN closely follows that of the Exa.TrkX collaboration, with minor deviations, as illustrated in Figure 4.
Figure 4: Schematic of the GNN architecture, highlighting: (1) hit and edge encodings, (2) $n$ message passing steps, and (3) subsequent edge and triplet classifications. The node and edge encoders, the node and edge networks and the edge and triplet classifiers are MLPs.
Initially, the hits $l$ are encoded from their normalised Cartesian coordinates $\vec{r}_{l}=(x_{l},y_{l},z_{l})$ into an $n_{h}$-dimensional space $\vec{h^{0}_{l}}\in\mathbb{R}^{n_{h}}$ using the node encoder (MLP). The concatenated node features $\vec{h^{0}_{i}}$ and $\vec{h^{0}_{j}}$ of edges $i\to j$ are input to the edge encoder (MLP), producing the edge encodings $e^{0}_{ij}\in\mathbb{R}^{n_{e}}$ in an $n_{e}$-dimensional space.
The hit and edge encodings are then iteratively updated over $n$ message passing steps. These steps allow the encodings to incorporate information from distant neighbours. During each message passing step $k\in\{0,\ldots,n-1\}$, a message $m^{k}_{l}$ is computed for each hit $l$ by aggregating the encodings of the edges connected to and from hit $l$. The sum aggregation is found to be sufficient for achieving acceptable performance. The message is computed as follows:
@@
Edge encodings are sufficient for both the edge MLP classifier and triplet MLP classifier. Therefore, the hit encodings are only utilised during the encoding and message passing steps to compute and update the edge encodings. The GNN is trained to classify both edges and triplets by minimizing the sum of the edge and triplet losses:
@@
We use sigmoid focal losses for the edge and triplet classification as it outperforms [30] the traditional binary cross-entropy loss. To ensure the GNN focuses on relevant triplets, edges with scores below 0.5 are discarded before triplet building and classification.
@@ 5.4 Track Building
Figure 6: Illustration of the four phases of track construction from the purified edge graph: (1) Connecting left and right elbows, (2) applying a WCC while omitting articulations with shared edges, (3) designating each residual link as a unique track, and (4) substituting edges with hits.
@@ 6 Physics Performance
Figure 7: Track-finding performance comparison for the search by triplet algorithm in Allen versus ETX4VELO for long particles as a function of the occupancy of the detector.
@@ 7 C++/CUDA Implementation and Throughput Measurement
The ETX4VELO pipeline is implemented in C++/CUDA using the Allen framework in order to measure its computational performance. Our implementation leverages Allen’s existing implementations for tasks such as memory management, event loading and dispatching, and VELO hit decoding. The classical Allen reconstruction pipeline is benchmarked dispatching 500 events to each of 16 CUDA streams and allocating 500 MB of GPU memory per stream. The ETX4VELO pipeline must adhere to these constraints to ensure optimal performance. Computational throughput is measured using the Nvidia GeForce RTX 2080Ti and GeForce RTX 3090 cards, with 50 repetitions of the pipeline in order to make the impact of I/O overhead on the throughput measurement negligible. The following steps of the pipeline have been implemented: (1) embedding network inference, (2) k-NN algorithm, (3) GNN inference up to the edge classifier, and (4) WCC algorithm. The implementation of track building from edge triplets is left for the future, but is not expected to have a major impact on the computational performance of the pipeline.
@@ 7.1 Network Inference
To infer the embedding network (MLP) and the GNN trained in PyTorch, inference engines are utilised. We have tried both ONNX Runtime (ORT) [34] with a CUDA backend and TensorRT (TRT) [35]. Both engines require the corresponding PyTorch networks to be exported to the ONNX open-source format. For TensorRT, memory allocation can be easily delegated to the Allen framework, making it well-suited for production and real-time inference scenarios. In contrast, ONNX Runtime does not readily support this feature. However, ONNX Runtime has the advantage of a CPU backend, making it easy to switch the pipeline from GPU to CPU when needed.

### v3

@@ 2 The LHCb detector and computing framework
The LHCb data acquisition can sustain a peak data rate of 5 TB/s which must be reduced to 10 GB/s before being written to long-term storage. This reduction is performed by a two-stage real-time processing system, the so-called trigger. The first stage is implemented on GPUs and is called Allen [^2] https://gitlab.cern.ch/lhcb/Allen [2]. Allen performs a partial event reconstruction, including tracking, and reduces the rate to between 70 and 200 GB/s using around 50 physics selections. The second stage is implemented on CPUs using the Moore [^3] https://gitlab.cern.ch/lhcb/Moore application based on the Gaudi [^4] https://gitlab.cern.ch/gaudi framework. Moore performs a full offline-quality reconstruction of the LHCb detector and reduces the rate to 10 GB/s using over two thousand physics selections.
Given this context, the Allen framework and LHCb offer an interesting use-case for developing Neural Network-based algorithms on GPUs, with a specific emphasis on high-throughput performance. In particular LHCb processes the full LHC collision rate using only around 320 GPU cards, leading to a minimum required throughput of around 80 kHz for the full processing pipeline on a single GPU. This in turn makes it essential to minimize overheads caused by data transfers between the host server and the GPU, which Allen achieves by processing batches of O(1000) events at a time. To the best of our knowledge this is the first time that a GNN tracking pipeline has been implemented using batched processing of this kind, or targeting these kinds of throughputs per GPU.
@@ 3 Track Topologies in the VELO
A VELO plane consists of four overlapping sensor layers, displaced along the $z$-axis, collaboratively covering the desired acceptance in the $\left(x,y\right)$-plane. The effective hit efficiency of the VELO is around 99% in simulation, so tracks must be allowed to skip a VELO plane during the reconstruction. A track can also leave more than one hit per plane, in most cases when it traverses overlapping sensor layers within a given plane. Material interactions frequently produce positron-electron pairs, resulting in two tracks that initially share hits before diverging. In fact 55% of electrons share hits with another particle, which motivates the need to take particular care when reconstructing them. Two tracks may also accidentally intersect, leading to a shared hit. The tracks that an effective VELO tracking algorithm should be capable of reconstructing are illustrated in Figure 2.
@@
4 Datasets
In order to have the information required for the training process of the machine learning models involved, such as the truth labels required for calculating the loss for a classification task, the pipeline is trained on simulated events. Similarly, for the evaluation of the accuracy of the algorithm, a simulation sample is necessary in order to have information about the particles produced in each event.
The results presented in this paper have been obtained using events produced with the full LHCb detector simulation in which $p$-$p$ collisions are generated using PYTHIA [23] and decayed using EvtGen [24]. The interaction of the generated particles with the detector, and its response, are implemented using the Geant4 toolkit [25] as described in [26]. We use simulated minimum bias samples corresponding to typical LHCb datataking conditions between 2022 and 2025, with an average of 5.3 inelastic $p$-$p$ collisions which produce at least one particle with momentum above 2 GeV in the LHCb detector acceptance per event. In our sample, there are, on average, 150 particles in the VELO acceptance, and 2,200 hits, in each event. In the simulation used for this paper around 15% of the hits are spilled over from prior events and therefore noise. The embedding network and GNN, introduced in Section 5, are trained on 700,000 events adhering to the selection criteria below.
  1. 1.
    Removal of non-linear particle tracks: The hits of the particle tracks that are not sufficiently linear due to multiple scattering (largely soft electrons) are removed. This is assessed by fitting a line to the particle hits and applying an upper limit to the average squared distance between the hits and the line. This criterion improves the physics performance during the training but also excludes 2.5% of the tracks reconstructible in the VELO.
  2. 2.
    Minimum number of VELO hits: A minimum of 500 genuine VELO hits is required.
  3. 3.
    Exclusion of tracks with insufficient hits: Tracks with fewer than 3 hits are excluded.
These criteria are not imposed on the test samples. Allen’s existing tracking algorithms, which are used by LHCb in 2024 datataking, are used as a reference for both the physics and computational performance of ETX4VELO, described in detail in Sections 6 and 7 respectively.
@@ 5 The ETX4VELO Pipeline
In general, a graph is a pair $G=(V,E)$, where $V$ is a finite set of vertices (or nodes), and $E$ is the set of connections (known as edges) between these nodes. In our problem, the hits left by charged particles in the VELO detector form a point cloud in 3-dimensional space. However, they can also be represented as a graph, in which the successive hits (or nodes) of each particle traversing the detector are connected to each other. This graph can be thought of as the “truth graph” that perfectly describes the VELO event. Conceptually, the end-goal of the ETX4VELO pipeline is to produce a graph that closely approximates this truth graph.
More specifically, the basic idea of our GNN pipeline is to build an initial graph of possible connections between hits in the detector, accurately classify these connections as correct (true/genuine) or incorrect (fake), and then, after discarding the fake ones, transform them into a set of track objects, containers of the hits of each track, which can then be understood and used by the rest of the algorithms in LHCb’s real-time pipeline. Since a graph with N nodes consisting of all possible node connections would have $C(N,2)=N(N-1)/2$ edges, and thus would be prohibitively large, a key challenge is to construct this initial graph in such a way that nearly all initial connections made are part of the final graph, while as many fake connections as possible are not. To illustrate this, for a typical VELO event of $N=2,200$ hits, a maximally connected graph would have $C(N,2)=5\times 10^{6}$ edges, while our method, described in the following paragraphs, creates on average 30,000 edges.
Our pipeline consists of 7 steps. We summarize the steps here, and then give further details on each step in dedicated subsections. Starting with the hits alone, the first two steps of the ETX4VELO pipeline construct a rough graph, $G^{\text{hit}}_{\text{rough}}$. Hits are first embedded into a Euclidean space using a Multi-Layer Perceptron (MLP). The MLP is trained to position hits that are likely to be connected by an edge close to each other in the embedding space. Subsequently, $k$-Nearest Neighbours ($k$-NN) algorithms are applied in the embedding space to collect edges that are most likely to be genuine. Nodes are considered at most two planes apart when identifying edge candidates.
@@
First, a GNN encodes each edge in $G^{\text{hit}}_{\text{rough}}$ ($i\to j$) into a high-dimensional vector $e_{i\to j}$, with dual training targets: edge classification and triplet classification. Second, the edge classifier network transforms these encodings into edge scores ranging from 0 (fake) to 1 (genuine). Edges with scores below $s_{\text{edge, min}}$ are discarded, resulting in the purified hit graph $G^{\text{hit}}_{\text{purified}}$. Third, the graph of edges $G^{\text{edge}}$ is built from the purified hit graph $G^{\text{hit}}_{\text{purified}}$. Finally, the triplet classifier network transforms the pairs of edge encodings (triplets) into triplet scores. Triplets scoring below $s_{\text{edge, min}}$ are discarded, resulting in the purified edge graph $G^{\text{edge}}_{\text{purified}}$.
@@ 5.1 Hit Embedding and Rough Graph Construction
To construct the graph $G^{\text{hit}}_{\text{rough}}$, one could connect each hit to all hits on the next two planes, allowing for the possibility of a missing plane due to pixel inefficiencies. However, this approach results in an excessive number of edges, thereby increasing the GNN’s inference time and memory usage. To enhance throughput, it is essential to minimise the graph size at this stage. We start by describing the operation of our method and then the training process and loss function in Eqs. (2) and (3).
@@
Using this trained embedding MLP, a hit on plane $p$ is connected to hits on the next two planes, $p+1$ and $p+2$, if they are within a squared distance of $d^{2}_{\text{max}}$. To avoid an excessive number of edges, a maximum of $k_{\text{max}}$ edges per node is imposed. Consequently, the rough graph $G^{\text{hit}}_{\text{rough}}$ is constructed by applying a $k_{\text{max}}$-NN algorithm on plane $p\in\left\llbracket 0,n_{\text{planes}}-1\right\rrbracket$ to the next two planes, $p+1$ and $p+2$, under a maximum squared distance of $d^{2}_{\text{max}}$. The k-NN implementation from faiss [28] is used for this purpose. The values of the hyperparameters $k_{\text{max}}$ and $d^{2}_{\text{max}}$ are determined post-training to balance the tradeoff between efficiency and clone rate, metrics described in Section 6. Here these values are chosen to be $k_{\text{max}}=50$ and $d^{2}_{\text{max}}=0.9$.
@@
The margin $m$, representing a squared distance threshold, is fixed at $m=1$. The parameter $w_{\text{genuine}}>1$ reduces the likelihood of excluding true edges within this margin, thus favouring the inclusion of genuine edges over the exclusion of false ones. The values $m$ and $w$, similary to $k_{\text{max}}$ and $d^{2}_{\text{max}}$, are chosen based on an elementary hyperparameter exploration.
@@ 5.2 Graph Neural Network and Classifiers
The Graph Neural Network (GNN) is employed to derive edge encodings used for both edge and triplet classification. The architecture of the GNN closely follows that of the Exa.TrkX collaboration, with minor deviations, as illustrated in Figure 4. The node and edge encoders, the node and edge networks and the edge and triplet classifiers are all MLPs, and should not be confused with the embedding network described in Section 5.1, which also happens to be an MLP.
Figure 4: Schematic of the GNN architecture, highlighting: (1) hit and edge encodings, (2) $n$ message passing steps, and (3) subsequent edge and triplet classifications. The node and edge encoders, the node and edge networks and the edge and triplet classifiers are all MLPs.
Initially, the hits $l$ are encoded from their normalised Cartesian coordinates $\vec{r}_{l}=(x_{l},y_{l},z_{l})$ into an $n_{h}$-dimensional space $\vec{h^{0}_{l}}\in\mathbb{R}^{n_{h}}$ using the node encoder. The concatenated node features $\vec{h^{0}_{i}}$ and $\vec{h^{0}_{j}}$ of edges $i\to j$ are input to the edge encoder, producing the edge encodings $e^{0}_{ij}\in\mathbb{R}^{n_{e}}$ in an $n_{e}$-dimensional space.
The hit and edge encodings are then iteratively updated over $n$ message passing steps. These steps allow the encodings to incorporate information from distant neighbours. During each message passing step $k\in\{0,\ldots,n-1\}$, a message $m^{k}_{l}$ is computed for each hit $l$ by aggregating the encodings of the edges connected to and from hit $l$. The message is computed as follows:
@@
Edge encodings are sufficient for both the edge classifier and triplet classifier. Therefore, the hit encodings are only utilised during the encoding and message passing steps to compute and update the edge encodings. The GNN is trained to classify both edges and triplets by minimizing the sum of the edge and triplet losses:
@@
We use sigmoid focal losses [30] for the edge and triplet classification since, for the purposes of our pipeline, it outperforms the traditional binary cross-entropy loss. To ensure the GNN focuses on relevant triplets, edges with scores below 0.5 are discarded before triplet building and classification.
@@ 5.4 Track Building
Figure 6: Illustration of the four phases of track construction from the purified edge graph using articulations and elbows defined in Section 5.3. (1) Connecting left and right elbows, (2) applying a WCC while omitting articulations with shared edges, (3) designating each residual link as a unique track, and (4) substituting edges with hits.
@@ 6 Physics Performance
Figure 7: Track-finding performance comparison for the search by triplet algorithm in Allen versus ETX4VELO for long particles, excluding electrons, as a function of the occupancy of the detector.
@@ 7 C++/CUDA Implementation and Throughput Measurement
The ETX4VELO pipeline is implemented in C++/CUDA using the Allen framework, the implementation of the first-level of the LHCb trigger on GPUs, in order to measure its computational performance. Our implementation leverages Allen’s existing implementations for tasks such as memory management, event loading and dispatching, and VELO hit decoding. The classical Allen reconstruction pipeline is benchmarked dispatching 500 events to each of 16 CUDA streams and allocating 500 MB of GPU memory per stream. The ETX4VELO pipeline must adhere to these constraints to ensure optimal performance. Computational throughput is measured using the Nvidia GeForce RTX 2080Ti and GeForce RTX 3090 cards, with 50 repetitions of the pipeline in order to make the impact of I/O overhead on the throughput measurement negligible. The following steps of the pipeline have been implemented: (1) embedding network inference, (2) k-NN algorithm, (3) GNN inference up to the edge classifier, and (4) WCC algorithm. The implementation of track building from edge triplets is left for the future, but is not expected to have a major impact on the computational performance of the pipeline.
@@ 7.1 Network Inference
To infer the embedding network and the GNN trained in PyTorch, inference engines are utilised. We have tried both ONNX Runtime (ORT) [34] with a CUDA backend and TensorRT (TRT) [35]. Both engines require the corresponding PyTorch networks to be exported to the ONNX open-source format. For TensorRT, memory allocation can be easily delegated to the Allen framework, making it well-suited for production and real-time inference scenarios. In contrast, ONNX Runtime does not readily support this feature. However, ONNX Runtime has the advantage of a CPU backend, making it easy to switch the pipeline from GPU to CPU when needed.